OpenStudy (anonymous):

Which fraction is NOT equivalent to 8/10?

OpenStudy (anonymous):

aight go homie

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

yes thats equal

OpenStudy (anonymous):

\[\frac{ 16 }{ 20 }\]

OpenStudy (anonymous):

to 8-10

OpenStudy (anonymous):

same

OpenStudy (anonymous):

-8/10 -8/-10

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cuz 2 negatives make a positive

OpenStudy (lukecrayonz):

Do the top number divided by the bottom :P 16/20=0.8. Or simplify, 16 and 20 both have 2 in common, divide by 2 and 8/10. and -8/10 is not equal to 0.8 as it is equal to -0.8

OpenStudy (anonymous):

ohh okay thanks guys :D

OpenStudy (anonymous):

give this guy a trophy :D

OpenStudy (anonymous):

ty.!

OpenStudy (anonymous):

Thanks for helping :D

OpenStudy (anonymous):

np